Factory Overhead Cost Variances

Perma Weave Textiles Corporation began January with a budget for 22,000 hours of production in the Weaving Department. The department has a full capacity of 29,000 hours under normal business conditions. The budgeted overhead at the planned volumes at the beginning of January was as follows:

Variable overhead $59,400
Fixed overhead 40,600
Total $100,000

The actual factory overhead was $101,200 for January. The actual fixed factory overhead was as budgeted. During January, the Weaving Department had standard hours at actual production volume of 23,000 hours. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Round your interim computations to the nearest cent, if required.

a. Determine the variable factory overhead controllable variance.
$   

b. Determine the fixed factory overhead volume variance.
$  
Answer #1
A. Variable factory overhead controllable variance.= actual variable factory over head -Budgeted variable factory overhead.
=$60600-$62100= $1500 F
Working Note
*Actual variable factory over head = total overhead - fixed over head
$101200 -$40600= $60600
Budgeted variable factory over head = Standard hours for actual output * variable overhead rate
= 23000 Hour X $2.70 = 62100
standard hours for actual output =23,000 hours.
variable overhead rate = budgeted overhead / budgeted hours
59400 /22,000 =$2,70
(b).
Fixed factory over head volume variance = (standard hours for 100% of normal capacity - standard hours for actual units produced )* Fixed factory overhead rate.
Fixed factory overhead volume variance = (29000 - 23,000)*$1.4= $8400 Unnfavourable
Working Note
standard hours for 100% of normal capacity = 29,000 hours.
standard hours for actual units produced = 23,000 hours.
Fixed factory overhead rate. = budeget fixed over head / capacity hours
= $40600 / 29,000= $1.4 per Hour